Information:

Using the money from their recent bond issue, Terry’s management has decided to declare an additional $562,500 dividend. The date of declaration is December 30, Year 3. The date of record will be January 15, Year 4, and the date of payment will be January 30, Year 4.

As an additional signal to the market, Terry’s management repurchased 205,000 shares of Terry’s common stock on December 15, Year 3 for $8.00 a share.

Terry’s management would like to know the effect of the sale on the following ratios:

*Current Ratio

*ROA

Assignment:

1. Calculate each of the two (2) ratios before you make any adjustments.

2. Make the appropriate journal entries, if any, to account for Terry’s extra dividend and stock repurchase (including any necessary changes to income tax expense).

3. Make any necessary changes to the financial statements.

4. Calculate the two (2) ratios after you make any adjustments.

Terry Co.
Balance Sheet
As of 12/31/Year 3
Year 3 Year 2
Assets
Current Assets
Cash $    5,554,969 $3,330,000
A/R $5,994,000 $5,661,000
Allowance for Bad Debts ($333,000) ($1,665,000)
Inventory $7,992,000 $9,324,000
Prepaid Insurance $499,500 $999,000
Prepaid Rent $832,500 $666,000
Total Current Assets $20,539,969 $18,315,000
Long-term Investments
Loans to other businesses $2,664,000 $2,664,000
Expansion Fund $2,843,560 $2,843,560
Total Long-term Investments $5,507,560 $5,507,560
PPE
Land $7,326,000 $4,662,000
Building $5,328,000 $5,328,000
Equipment $18,648,000 $8,658,000
Accumulated Depreciation ($8,658,000) ($6,660,000)
Total PPE $22,644,000 $11,988,000
Intangible Assets
Patents, net $999,000 $999,000
Total Assets $49,690,529 $36,809,560
Liabilities and Stockholders' Equity
Current Liabilities
Accounts Payable $3,044,570 $3,996,000
Income Tax Payable $    2,322,880 $666,000
Interest Payable $        30,000 $0
Unearned Revenue $1,198,800 $999,000
Wages Payable $666,000 $832,500
Current Portion of Loan Payable $333,000 $333,000
Total Current Liabilities $7,595,250 $6,826,500
Long-term Debt
Loan Payable $3,663,000 $3,996,000
Notes Payable $9,324,000 $5,328,000
Bonds Payable $    1,561,447 $0
Total Long-term Debt $14,548,447 $9,324,000
Total Liabilities $22,143,697 $16,150,500
Stockholders' Equity
Common stock   $2,660,000 $2,660,000
   ($1 par, 4,655,000 authorized, 2,660,000 outstanding)
Additional Paid-In capital $1,998,000 $1,998,000
Retained Earnings $23,633,803 $16,746,032
Accumulated OCI ($744,972) ($744,972)
Total Stockholders' Equity $27,546,831 $20,659,060
Total Liabilities and Stockholder's Equity $49,690,528 $36,809,560
Answer #1

1. Current ratio before making any adjustments = current assets (before adjustments )/ current liability (before adjustment)

= 20539969/7595250

= 2.70

ROA (before adjustment ) = net earnings / total assets

= Please note that for calculating ROA , I need statement of profitability and loss to get the figure of Net earnings, since profit and loss statement is not given, it is not possible to calculate.

2. Journal entries would be as follows (IFRS)

Date general journal debit Credit
Dec 15 year 3

treasury stock

Cash

(Being 2050000 shares repurchase @ $8 per share

1640000

1640000

Dec 31 year 3

retained earnings

Dividend payable

(To dividend declared )

562500

562500

3 . First of all Cash for year 3 will get reduced as 5554969-1640000 = $3914969.

Current assets total = 20539969-1640000 = $18899969

2. Current liabilities will also include dividend payable amounting to $562500.

Total of current liabilities = 7595250+562500 = $8157750

3 . Stockholders equity section would be as follows :-

Common stock 2660000
Additional paid in capital 1998000
Total paid in capital 4658000
Retained earnings (23633803-562500) 23071303
OCI (744972)
Subtotal $26984331
Less : treasury stock (205000 shares @$8 per share) (1640000)
Total stockholder equity $25344331

3 . Current ratio after adjustment = current assets (after adjustment )/ current liabilities (after adjustment )

= 18899969/8157750 = 2.32.

Again for ROA net earning is required